Could I ask for your help?
What will the last payment be to pay off a $1,250 computer plus 8% sales tax with monthly payments of $120?
$10
$11
$12
$20
$30
Four people in a room. Each shakes hands with each of the others once. How many handshakes are there?
6
10
12
16
22

Answer by KnightOwlTutor(293)

What will the last payment be to pay off a $1,250 computer plus 8% sales tax with monthly payments of $120?
We need to calculate the sales tax and add this amount to the cost of the computer.
convert 8% to decimal form =8/100 =.08
1250*.08=100
1250+100=1350
To find out how many payments of 120 can be made
120X=1350
divide both sides by 120
X=11.25 payments of $120
since .25 is a fraction we can calculate the actual $ amount by multiplying 120*.25=$30
Thus the last payment is $30
Four people in a room. Each shakes hands with each of the others once. How many handshakes are there?

Let's make a chart
we assign a # to each of the 4 persons
1,2,3,4
1 shakes hands with 2,3,4
2 shakes hands with 3,4
3 shakes hands with 4
add up each case 3+2+1=6 total # of handshakes